OpenStudy (anonymous):

in a ||gm if A=115° then find B C D

OpenStudy (anonymous):

1) a=c = 115 , a+c = 2*115 2) b=d = (360-2*115)/2 reason 1) opposite angles of a parallelogram are equal 2) sum of angles of a quadrilateral is 360 degree

OpenStudy (bradely):

\[A=C=115 and B=D=x\] Total angle 360 A+B+C+D=360 115+x+115+x=360 2x+230=360 2x=130 x=65